# Automatically Interpreting Millions of Features in Large Language Models

### v2

@@ A.2 Explainer Prompt
the text latents that are common in the examples, and what
@@
We add to the previous prompt the following if we want to do COT: {mdframed}
@@
2. Write down general shared latents of the text examples.
@@ 3. Formulate a hypothesis and write down the final interpretation
One of the few shot examples of how examples are displayed to the model. {mdframed}
@@
If COT is used, an explicit example of using COT is demonstrated. {mdframed}
@@
Afterwards an interpretation is added to the example {mdframed}
@@
A.3 Examples of activating contexts and interpretations
As discussed in the main text, the interpretations found for a given latent can be very different depending on the way to sample the activating contexts shown to the explainer model, see Figure A2. This has its advantages and disadvantages.
@@ A.3 Examples of activating contexts and interpretations
On the other hand, if we look at the second example, sampling from the whole distribution examples may sometimes confuse the explainer model - “Tokens often precede or succeed prepositions, articles, and words that signal possession or quantity, frequently indicating a relationship between objects or actions. of latent”, which might not see the pattern that is more clear in the top activating examples -“Descriptions of food or events where food is involved, often mentioning leftovers, and sometimes mentioning the act of eating, serving, or storing food, as well as the amount of food, or the pleasure or satisfaction derived from it.” Here we see very poor scores on the interpretation from randomly sampled examples, 0.54 accuracy in detection and 0.57 accuracy in fuzzing, while the interpretation generated from the top examples has 0.89 accuracy both in detection and fuzzing,
Figure A2: Activating contexts of latent 209 (top) and 293 (bottom), from layers 8 and 32, respectively, of the 131k latent SAE trained on the residual stream of Gemma 2 9b. The shown examples are similar to the ones given to the explainer model to come up with interpretations.
A.4 Different scoring methods
A.4.1 Detection details
The prompt used for detection scoring is the following: {mdframed}
@@ A.4.1 Detection details
You will be given a certain latent of text, such as
@@
is to determine which examples possess the latent.
@@
Together with the prompt, there are several few shot examples like the following: {mdframed}
@@
latent interpretation: Words related to American football
@@
In this example the $<$user_prompt$>$ and the $<$assistant_response$>$ tags are substituted with the correct instruct format used by the scorer model. In detection scoring, 5 shuffled examples are shown to the model at the same time.
A.4.2 Fuzzing details
The prompt used for fuzzing scoring is the following: {mdframed}
@@ A.4.2 Fuzzing details
You will be given a certain latent of text, such as ‘‘male
@@
be given a few examples of text that contain this latent.
@@
latent are between tokens << and >>.
@@
Followed by few-shot examples of the kind: {mdframed}
@@
latent interpretation: Words related to American football
@@
In this example, the $<$user_prompt$>$ and the $<$assistant_response$>$ tags are substituted with the correct instruct format used by the scorer model. In fuzzing scoring, 5 shuffled examples are shown to the model at the same time.
A.4.3 Surprisal details
In surprisal scoring, the cross-entropy loss of the scorer model is computed over the tokens of an example. For each interpretation, this loss is computed with the interpretation and with a default interpretation - “Various unrelated sentences,” where the examples can either be activating context or non-activating contexts. In our first approach, Llama 3.1 70b base was used. The prompt starts with few shot examples like: {mdframed}
The following is a description of a certain latent of text
and a list of examples that contain the latent.
@@ A.4.3 Surprisal details
 the leading latent which we ought  "
@@
With the pairs of losses with the interpretation and with the default interpretation, it is possible to compute the decrease in loss caused by having access to the interpretation. It is expected that in activating contexts this difference will be greater than in non-activating examples, and the score of the interpretation is given by the AUC computed using this loss as a proxy for activating and non-activating labels.
A.4.4 Embedding details
@@
A set of activating and randomly selected non-activating examples are embedded using the scorer model. Then a “query” instruction is embedded: {mdframed}
@@
A.4.5 Adversarial examples
Most latents are active in less than 0.1% of the full dataset used to collect the activations, making random non-activating contexts very diverse. Randomly sampling non-activating examples cannot be used to determine whether a latent fires in a token in a specific context or on that token in general, as it is unlikely that that token randomly occurs in each non-activating example. As SAEs are scaled and latents become sparser and more specific, techniques that overly rely on activating contexts will have more imprecise results Gao et al. (2024).
Motivated by the phenomenon of latent splitting, we could use “similar” latents to test whether interpretations are precise enough to distinguish between similar contexts. A potential approach is using cosine similarity between decoder directions of latents to find counterexamples for an interpretation. Some works (Juang et al., 2024; Macgrath, 2024) have shown that a significant fraction of current latent interpretations can’t be used to distinguish between latents with high similarity.
A.4.6 Correlation between scores
We compute the correlation between these different scores and simulation 800 different latent interpretations, of the 131k latent SAE trained on the residual stream of Gemma 2 9b spread across 4 different layers. We use 100 activating example, 10 from each of 10 different activating quantiles, and 100 non-activating examples to compute these scores and use Llama 3.1 70b Instruct as the scorer model for simulation, fuzzing and detections, while we use Llama 3.1 70b Base for surprisal and Stella-400M as the embedding model.
Interestingly, we also find that the correlation between simulation scores and other scores also increases when the simulation scorer is Claude Sonnet 3.5. These estimations are on a smaller number of latents(100), due to the high cost of simulation using Claude Sonnet, but on this set, the Spearman correlation between fuzzing and simulation increases from 0.69 to 0.75, between detection and simulation increases from 0.31 to 0.38, between embedding and simulation from 0.14 to 0.21 and from surprisal to simulation from 0.02 to 0.10. The baseline numbers for the correlation of simulation scores are lower on this set, so we expect that had we done Claude scores on a larger set of explanations, the correlations between the scores would be higher.
Table A1: Spearman correlation, computed over scores of 800 different latent interpretations.
@@ A.4.6 Correlation between scores
Table A2: Pearson correlation computed over 800 different latent scores
@@
A.4.7 Disagreement between scores
@@
For instance, feature 2 of layer 8 of the 131k latent SAE model trained on the residual stream of Gemma 2 9b has a fuzzing score of 0.9 but a detection score of only 0.43. The interpretation given by our pipeline is ”Verbs that link a subject to additional information, often in a formal or descriptive tone.” Most of the activations of this latent are on sentences like ”This $<$is needed$>$ because Magento (…)” or ”For instance, this$<$ is$>$ technically correct syntax”, where <>absent<>< > represent active tokens. In both these examples, the model correctly identifies the highlighted tokens as active during fuzzing, but incorrectly identifies the context as non-active during detection. On the other hand, detection incorrectly identifies sentences like ”pay for things that would prevent larger issues down the road is better in the long run.” or ”Neuroscientist Jack Gallant calls the research a technologic tour de force and says the ultimate decoder would provide vivid” as active. This explanation is too vague on the activating context, leading to a low detection score, but specific enough in the types of tokens that are active, leading to a high fuzz score.
Feature 281 of the same layer has a fuzzing score of 0.97, but a simulation score of only 0.19. Its interpretation is ”URLs or hyperlinks containing query parameters, indicating a request for specific data or actions on a web page.” The simulation score is lower for than the fuzzing score because the model has to decide which parts of the URL the latent is active on, while the fuzzing scorer either is shown links that are highlighted which it will say are active or non-links that are highlighted and are, most likely, not active.
@@
A.5 SAE models used
Throughout this work, we used different SAEs trained on Gemma. The 16k latent ones trained on the MLP have the following average L0 norms per layer: {mdframed} 0:50, 1:56, 2:33, 3:55, 4:66, 5:46, 6:46, 7:47, 8:55, 9:40, 10:49, 11:34, 12:42, 13:40, 14:41, 15:45, 16:37, 17:41, 18:36, 19:38, 20:41, 21:34, 22:34, 23:73, 24:32, 25:72, 26:57, 27:52, 28:50, 29:49, 30:51, 31:43, 32:44, 33:48, 34:47, 35:46, 36:47, 37:53, 38:45, 39:43, 40:37, 41:58 The 16k latent ones trained on the residual stream have the following average L0 norms: {mdframed} 0:35, 1:69, 2:67, 3:37, 4:37, 5:37, 6:47, 7:46, 8:51, 9:51, 10:57, 11:32, 12:33, 13:34, 14:35, 15:34, 16:39, 17:38, 18:37, 19:35, 20: 36, 21:36, 22: 35, 23: 35, 24: 34, 25: 34, 26: 35, 27:36, 28: 37, 29:38, 30:37, 31:35, 32: 34, 33:34, 34:34, 35:34, 36:34, 37:34, 38:34, 39:34, 40:32, 41:52 The 131k latent ones trained in the residual steam have the following average L0 norms: {mdframed} 0:30, 1:33, 2:36, 3:46, 4:51, 5:51, 6:66, 7:38, 8:41, 9:42, 10:47, 11:49, 12:52, 13:30, 14:56, 15:55, 16:35, 17:35, 18:34, 19:32, 20:34, 21:33, 22:32, 23:32, 24: 55, 25:54, 26:32, 27:33, 28: 32, 29:33, 30:32, 31:52, 32: 51, 33:51, 34:51, 35:51, 36: 51, 37:53, 38:53, 39:54, 40: 49, 41:45
We also used SAEs with 65k latents, trained on the residual stream and the MLP of Llama 8b.
A.6 Factors that influence the explainability
A.6.1 Dependence on dataset
Even though a significant portion of SAE latents are less active when using RPJv2 instead of the Pile, we find that the latents that are active are generally interpretable to the same degree. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.6.1 Dependence on dataset
A.6.2 Dependence on chain of thought and activation information
We find that COT slightly increases the scores of the interpretations found, and that it significantly slows down the rate at which one can produce interpretations. Giving the explainer model, the activations associated with each token seem to slightly increase the scores of the interpretations generated. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.6.2 Dependence on chain of thought and activation information
A.6.3 Dependence on context length
The context length of the shown examples did not significantly change the scores obtained for the interpretations. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ A.6.3 Dependence on context length
A.6.4 Dependence on the origin of examples
Sampling the examples from the whole distribution of activations, or sampling a fixed number of examples over different activation quantiles, significantly improved the scores, compared with sampling only from the top examples. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ A.6.4 Dependence on the origin of examples
A.6.5 Dependence on the number of examples
The number of examples shown to the model seems to saturate, at least with the explainer model we used. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.6.5 Dependence on the number of examples
A.6.6 Explainability across layers
We find that, in the case of the 131k latent SAE trained on the residual stream, the earliest layers have lower scores than the later layers. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
Figure A5: Accuracy on fuzzing and detection scoring Dots correspond to the median score over c.a. 300 latent interpretations, and colored region denotes the interquartile range.
A.6.7 Dependence on size of explainer models
Having a larger explainer model leads to better interpretations, but our results suggest that this benefit tends to saturate, as we find that the scores of explanations given by Claude or Llama 70b are very similar. We only did human interpretation of 100 latents. Simulations scores are also only computed over 100 latents.
@@ A.6.7 Dependence on size of explainer models
A.6.8 Dependence on size of explainer models
@@
A.6.9 Dependence on SAE size and location
We compare the interpretability of SAEs with different number of latents, trained on the same model, with the neurons of that model. We find that the SAE with the highest number of latents to have the highest scores in the case of the Gemma 2 9b SAEs and that SAEs trained on the residual stream have higher scores for both Gemma 2 9b and Llama 3.1 8b.

### v3

@@ Appendix B Explainer Prompt
the text features that are common in the examples, and what
@@
We add to the previous prompt the following if we want to do COT:
@@
2. Write down general shared features of the text examples.
@@ 3. Formulate a hypothesis and write down the final interpretation
One of the few shot examples of how examples are displayed to the model.
@@
If COT is used, an explicit example of using COT is demonstrated.
@@
Afterwards an interpretation is added to the example
@@
Appendix C Examples of activating contexts and interpretations
As discussed in the main text, the interpretations found for a given feature can be very different depending on the way to sample the activating contexts shown to the explainer model, see Figure A2. This has its advantages and disadvantages.
@@ Appendix C Examples of activating contexts and interpretations
On the other hand, if we look at the second example, sampling from the whole distribution examples may sometimes confuse the explainer model - “Tokens often precede or succeed prepositions, articles, and words that signal possession or quantity, frequently indicating a relationship between objects or actions. of feature”, which might not see the pattern that is more clear in the top activating examples -“Descriptions of food or events where food is involved, often mentioning leftovers, and sometimes mentioning the act of eating, serving, or storing food, as well as the amount of food, or the pleasure or satisfaction derived from it.” Here we see very poor scores on the interpretation from randomly sampled examples, 0.54 accuracy in detection and 0.57 accuracy in fuzzing, while the interpretation generated from the top examples has 0.89 accuracy both in detection and fuzzing,
Figure A2: Activating contexts of feature 209 (top) and 293 (bottom), from layers 8 and 32, respectively, of the 131k feature SAE trained on the residual stream of Gemma 2 9b. The shown examples are similar to the ones given to the explainer model to come up with interpretations.
Appendix D Different scoring methods
D.1 Detection details
The prompt used for detection scoring is the following:
@@ D.1 Detection details
You will be given a certain feature of text, such as
@@
is to determine which examples possess the feature.
@@
Together with the prompt, there are several few shot examples like the following:
@@
feature interpretation: Words related to American football
@@
In this example the <<user_prompt>> and the <<assistant_response>> tags are substituted with the correct instruct format used by the scorer model. In detection scoring, 5 shuffled examples are shown to the model at the same time.
D.2 Fuzzing details
The prompt used for fuzzing scoring is the following:
@@ D.2 Fuzzing details
You will be given a certain feature of text, such as ‘‘male
@@
be given a few examples of text that contain this feature.
@@
feature are between tokens << and >>.
@@
Followed by few-shot examples of the kind:
@@
feature interpretation: Words related to American football
@@
In this example, the <<user_prompt>> and the <<assistant_response>> tags are substituted with the correct instruct format used by the scorer model. In fuzzing scoring, 5 shuffled examples are shown to the model at the same time.
D.3 Surprisal details
In surprisal scoring, the cross-entropy loss of the scorer model is computed over the tokens of an example. For each interpretation, this loss is computed with the interpretation and with a default interpretation - “Various unrelated sentences,” where the examples can either be activating context or non-activating contexts. In our first approach, Llama 3.1 70b base was used. The prompt starts with few shot examples like:
The following is a description of a certain feature of text
and a list of examples that contain the feature.
@@ D.3 Surprisal details
 the leading feature which we ought  "
@@
With the pairs of losses with the interpretation and with the default interpretation, it is possible to compute the decrease in loss caused by having access to the interpretation. It is expected that in activating contexts this difference will be greater than in non-activating examples, and the score of the interpretation is the AUC computed using this loss as a proxy for activating and non-activating labels.
D.4 Embedding details
@@
A set of activating and randomly selected non-activating examples are embedded using the scorer model. Then a “query” instruction is embedded:
@@
D.5 Adversarial examples
Most features are active in less than 0.1% of the full dataset used to collect the activations, making random non-activating contexts very diverse. Randomly sampling non-activating examples cannot be used to determine whether a feature fires in a token in a specific context or on that token in general, as it is unlikely that that token randomly occurs in each non-activating example. As SAEs are scaled and features become sparser and more specific, techniques that overly rely on activating contexts will have more imprecise results (Gao et al., 2024).
Motivated by the phenomenon of feature splitting, we could use “similar” features to test whether interpretations are precise enough to distinguish between similar contexts. A potential approach is using cosine similarity between decoder directions of features to find counterexamples for an interpretation. Some works (Juang et al., 2024; Macgrath, 2024) have shown that a significant fraction of current feature interpretations can’t be used to distinguish between features with high similarity.
D.6 Correlation between scores
We compute the correlation between these different scores and simulation 800 different feature interpretations, of the 131k feature SAE trained on the residual stream of Gemma 2 9b spread across 4 different layers. We use 100 activating example, 10 from each of 10 different activating quantiles, and 100 non-activating examples to compute these scores and use Llama 3.1 70b Instruct as the scorer model for simulation, fuzzing and detections, while we use Llama 3.1 70b Base for surprisal and Stella-400M as the embedding model.
Interestingly, we also find that the correlation between simulation scores and other scores also increases when the simulation scorer is Claude Sonnet 3.5. These estimations are on a smaller number of features(100), due to the high cost of simulation using Claude Sonnet, but on this set, the Spearman correlation between fuzzing and simulation increases from 0.69 to 0.75, between detection and simulation increases from 0.31 to 0.38, between embedding and simulation from 0.14 to 0.21 and from surprisal to simulation from 0.02 to 0.10. The baseline numbers for the correlation of simulation scores are lower on this set, so we expect that had we done Claude scores on a larger set of explanations, the correlations between the scores would be higher.
Table A1: Spearman correlation, computed over scores of 800 different feature interpretations.
@@ D.6 Correlation between scores
Table A2: Pearson correlation computed over 800 different feature scores
@@
D.7 Disagreement between scores
@@
For instance, feature 2 of layer 8 of the 131k feature SAE model trained on the residual stream of Gemma 2 9b has a fuzzing score of 0.9 but a detection score of only 0.43. The interpretation given by our pipeline is ”Verbs that link a subject to additional information, often in a formal or descriptive tone.” Most of the activations of this feature are on sentences like ”This <<is needed>> because Magento (…)” or ”For instance, this<< is>> technically correct syntax”, where <><> represent active tokens. In both these examples, the model correctly identifies the highlighted tokens as active during fuzzing, but incorrectly identifies the context as non-active during detection. On the other hand, detection incorrectly identifies sentences like ”pay for things that would prevent larger issues down the road is better in the long run.” or ”Neuroscientist Jack Gallant calls the research a technologic tour de force and says the ultimate decoder would provide vivid” as active. This explanation is too vague on the activating context, leading to a low detection score, but specific enough in the types of tokens that are active, leading to a high fuzz score.
Feature 281 of the same layer has a fuzzing score of 0.97, but a simulation score of only 0.19. Its interpretation is ”URLs or hyperlinks containing query parameters, indicating a request for specific data or actions on a web page.” The simulation score is lower for than the fuzzing score because the model has to decide which parts of the URL the feature is active on, while the fuzzing scorer either is shown links that are highlighted which it will say are active or non-links that are highlighted and are, most likely, not active.
@@
Appendix E SAE models used
Throughout this work, we used different SAEs trained on Gemma. The 16k feature ones trained on the MLP have the following average L0 norms per layer: 0:50, 1:56, 2:33, 3:55, 4:66, 5:46, 6:46, 7:47, 8:55, 9:40, 10:49, 11:34, 12:42, 13:40, 14:41, 15:45, 16:37, 17:41, 18:36, 19:38, 20:41, 21:34, 22:34, 23:73, 24:32, 25:72, 26:57, 27:52, 28:50, 29:49, 30:51, 31:43, 32:44, 33:48, 34:47, 35:46, 36:47, 37:53, 38:45, 39:43, 40:37, 41:58 The 16k feature ones trained on the residual stream have the following average L0 norms: 0:35, 1:69, 2:67, 3:37, 4:37, 5:37, 6:47, 7:46, 8:51, 9:51, 10:57, 11:32, 12:33, 13:34, 14:35, 15:34, 16:39, 17:38, 18:37, 19:35, 20: 36, 21:36, 22: 35, 23: 35, 24: 34, 25: 34, 26: 35, 27:36, 28: 37, 29:38, 30:37, 31:35, 32: 34, 33:34, 34:34, 35:34, 36:34, 37:34, 38:34, 39:34, 40:32, 41:52 The 131k feature ones trained in the residual steam have the following average L0 norms: 0:30, 1:33, 2:36, 3:46, 4:51, 5:51, 6:66, 7:38, 8:41, 9:42, 10:47, 11:49, 12:52, 13:30, 14:56, 15:55, 16:35, 17:35, 18:34, 19:32, 20:34, 21:33, 22:32, 23:32, 24: 55, 25:54, 26:32, 27:33, 28: 32, 29:33, 30:32, 31:52, 32: 51, 33:51, 34:51, 35:51, 36: 51, 37:53, 38:53, 39:54, 40: 49, 41:45
We also used SAEs with 65k features, trained on the residual stream and the MLP of Llama 8b.
Appendix F Factors that influence the explainability
F.1 Dependence on dataset
Even though a significant portion of SAE features are less active when using RPJv2 instead of the Pile, we find that the features that are active are generally interpretable to the same degree. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ F.1 Dependence on dataset
F.1.1 Dependence on chain of thought and activation information
We find that COT slightly increases the scores of the interpretations found, and that it significantly slows down the rate at which one can produce interpretations. Giving the explainer model, the activations associated with each token seem to slightly increase the scores of the interpretations generated. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ F.1.1 Dependence on chain of thought and activation information
F.2 Dependence on context length
The context length of the shown examples did not significantly change the scores obtained for the interpretations. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ F.2 Dependence on context length
F.3 Dependence on the origin of examples
Sampling the examples from the whole distribution of activations, or sampling a fixed number of examples over different activation quantiles, significantly improved the scores, compared with sampling only from the top examples. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ F.3 Dependence on the origin of examples
F.4 Dependence on the number of examples
The number of examples shown to the model seems to saturate, at least with the explainer model we used. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ F.4 Dependence on the number of examples
F.5 Explainability across layers
We find that, in the case of the 131k feature SAE trained on the residual stream, the earliest layers have lower scores than the later layers. These evaluations were done using the 131k feature SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
Figure A5: Accuracy on fuzzing and detection scoring Dots correspond to the median score over c.a. 300 feature interpretations, and colored region denotes the interquartile range.
F.6 Dependence on size of explainer models
Having a larger explainer model leads to better interpretations, but our results suggest that this benefit tends to saturate, as we find that the scores of explanations given by Claude or Llama 70b are very similar. We only did human interpretation of 100 features. Simulations scores are also only computed over 100 features.
@@ F.6 Dependence on size of explainer models
F.7 Dependence on size of explainer models
@@
F.8 Dependence on SAE size and location
We compare the interpretability of SAEs with different number of features, trained on the same model, with the neurons of that model. We find that the SAE with the highest number of features to have the highest scores in the case of the Gemma 2 9b SAEs and that SAEs trained on the residual stream have higher scores for both Gemma 2 9b and Llama 3.1 8b.